OpenStudy (anonymous):

7log4^(2x-8)= log4^4x+4

OpenStudy (anonymous):

Just checking. Should that 4x+4 at the end be in parentheses?

OpenStudy (anonymous):

No it's open.

OpenStudy (anonymous):

Is it \[\log 4^{4x+4}\] or \[\log 4^{4x} + 4\]

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

The log identity to use here is \[a \log b = \log b^a\]

OpenStudy (anonymous):

You've been looking at this question a while now Marina. Feel free to chime in lol. All help is welcomed

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

Thanks to you both. May I ask how you arrived at that answer Marina?

OpenStudy (anonymous):

By using the log identity dmancine gave you. In addition, it's very helpful if you will use as many parenthesis as needed to avoid any confusion.

OpenStudy (anonymous):

7(2x-8)log4=(4x+4)log4

OpenStudy (anonymous):

14x-56=4x+4 10x=60 x=6

OpenStudy (anonymous):

Just to get clarification for the future, should the bases be different, do you use the same log identity?

OpenStudy (anonymous):

If you really mean the base of the log, then you can use the identity\[\log _{a} b = (\log b) / (\log a)\]

OpenStudy (anonymous):

Ok I figured that. Just wanted to make sure.

OpenStudy (anonymous):

I am glad that both of us were helful :-)